y = -1/2x + 1; (4, 2) how do i solve for the parallel for (4,2)
Answer by oscargut(2103)

parallel must have the same slope then:
y = -1/2 x + b
Since (4,2) is in the line:
2 = (-1/2)4+b
2 = -2+b
b = 4
Answer: y = -1/2 x +4
You can ask me more at: mthman@gmail.com
Thanks